OpenStudy (jhannybean):

Calculus 3: Convert \(3x^2-8xy+3y^2+4z^2=7\) using cylindrical coordinates, I just need something checked, not solved.

OpenStudy (jhannybean):

I'll just work it out here since it's harder to see on Skype I guess? lol

OpenStudy (jhannybean):

Since you prefer being tagged @perl

OpenStudy (perl):

3 ( x^2 + y^2 ) , thats a nice expression

OpenStudy (jhannybean):

\[3x^2 -8xy +3y^2 +4z^2 = 7\]\[3(rcos(\theta))^2 -8(r\cos(\theta))(r\sin(\theta)) +3(rsin(\theta))^2 +4z^2 = 7\]\[3r^2\cos^2(\theta) -8r^2\cos(\theta)\sin(\theta)+3r^2sin^2(\theta) + 4z^2 = 7\]\[3r^2(cos^2(\theta) +sin^2(\theta)) \color{red}{-8r^2cos(\theta)sin(\theta)}+4z^2=7\]\[3r^2 \color{red}{-8r^2\cos(\theta)sin(\theta)} +4z^2 = 7\]

OpenStudy (perl):

3(x^2 + y^2) , that becomes 3*r^2

OpenStudy (jhannybean):

you're saying... and i'm repeating this to understand it better...\[\sin(2\theta) = 2sin(\theta)cos(\theta)\]\[-4\cdot sin(2\theta) = -4 \cdot 2sin(\theta)cos(\theta))\]

OpenStudy (jhannybean):

Yeah, I got that part @perl haha, I was confused about the red portion.

OpenStudy (perl):

wow youre really good at Latex

OpenStudy (perl):

you can even do colors

OpenStudy (anonymous):

I have no idea how she highlights with colors, lol.

OpenStudy (jhannybean):

`\color{red}{@Concentrationalizing }`

OpenStudy (perl):

cylindrical coordinates are ( r , theta, z )

OpenStudy (jhannybean):

\[\color{red}{Concentrationalizing}\]

OpenStudy (perl):

so if you have an equation in terms of r, theta, z , you are done

OpenStudy (jhannybean):

Yep those are cylindrical coordinates.

OpenStudy (anonymous):

FAIL!

OpenStudy (jhannybean):

i posted this question up because he posted up a neat trick in skype, and I just wanted to see if I could replicate it here, haha. I had forgotten that \(\sin(2\theta) = 2sin(\theta)cos(\theta)\)

OpenStudy (perl):

is that the final answer, can we simplify it more

OpenStudy (jhannybean):

Possibly simplify it even more. What is fail? @Concentrationalizing >:((

OpenStudy (perl):

3r^2 - 4r^2 sin(2theta) = 7 - 4z^2 (3 - 4sin(2theta) r^2 = 7 - 4z^2

OpenStudy (jhannybean):

you have to identify what red is, which is a `color` :P

OpenStudy (jhannybean):

\[3r^2 \color{red}{-8r^2\cos(\theta)sin(\theta)} +4z^2 = 7\]\[3r^2 \color{red}{-4sin(2\theta)}+4z^2 =7\]\[4z^2 = 7 -3r^2+4\sin(2\theta)\]

OpenStudy (jhannybean):

Then we solve for z.....right? O_o

OpenStudy (jhannybean):

Wait - @perl why are you subtracting \(4z^2\)?

OpenStudy (perl):

you can do it that way

OpenStudy (perl):

thats cleaner

OpenStudy (jhannybean):

So that means we were solving for r^2 this whole time?!

OpenStudy (jhannybean):

Or we're trying to parallel the equation of a circle.. hm.

OpenStudy (perl):

nope, i dont think we have to solve for r, or z , or theta, in particular

OpenStudy (perl):

as long as you have an equation in terms of r, z, theta only, youre good to go

OpenStudy (perl):

as long as you dont have any x or y

OpenStudy (perl):

my attach file button is broken but

OpenStudy (perl):

question 1

OpenStudy (perl):

he didn't solve for r or z

OpenStudy (jhannybean):

I see

OpenStudy (jhannybean):

\[3r^2 \color{red}{-8r^2\cos(\theta)sin(\theta)} +4z^2 = 7\]\[r^2 (3-4\sin(2\theta)) +4z^2 = 7\]\[r^2 = \frac{7-4z^2}{3-4sin(2\theta)}\]

OpenStudy (jhannybean):

\begin{array}{r|lll} 6&1&-7&6\\ \\ &\hline 1 \end{array}